OpenStudy (anonymous):

If f(x) = x^2 - 1, find f(3).

OpenStudy (lgbasallote):

simply substitute 3 into x f(x) = x^2 - 1 f(3) = (3)^2 - 1 does that help?

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome ^_^